OpenStudy (itsmichelle29):

A motorcycle cost $12,000 when it was purchased. The value of a motorcycle decreases by 6% each year. Find the rate of decay each month and select the correct answer below. −0.005143% −0.5143% −0.005% −0.5%

OpenStudy (itsmichelle29):

i medal and fan I need help with two questions

OpenStudy (anonymous):

12,000 doesnt matter. Since they are asking for each year that means divide by 12. They already give you a rate for decay they just want a conversion. So take 6% and convert it into decimal form by dividing it by 100 \[\frac{ 6 }{ 100 } = .06\] Now since it is per month divide by 12 \[\frac{ .06 }{ 12 } = .005\] Now since all the answers are in percentages multiply .005 by 100 to put it into that form. \[\left( .005 \right)\left( 100 \right) = .5 percent\]